OpenStudy (anonymous):

Rupert drove home at an average rate of 58 miles per hour. Find his rate in feet per seconds.

OpenStudy (campbell_st):

how many feet in 58 miles..?

OpenStudy (campbell_st):

when you get the answer divide by 60 to get the speed as feet per minute then divide by 60 again to get feet per second hope it helps

OpenStudy (tkhunny):

Multiply by one (1) \(\dfrac{58\;mile}{hour}\cdot\dfrac{5280\;ft}{1\;mile}\cdot\dfrac{1\;hour}{3600\;sec}\)

OpenStudy (anonymous):

\[58 \frac{miles}{hour}\] since 1mile = 5280 feet and 1hour = 60² seconds you just replace one for the other since they are equal: \[58 \frac{mile}{hour} = 58 \frac{5280 feet}{60^{2} seconds} = \frac{58*5280}{3600}\frac{feet}{seconds}\] which is 85.06ft/s